# Avalanche Dynamics and the Effect of Straining in Dislocation Systems with Quenched Disorder

### v1

@@ III.5 Yield stresses and local yield stresses
After studying the correlations and the avalanche behavior now we turn to one of the main questions if this work: how are the local yield stresses affected by the avalanches and the deformation of the model sample. Firstly, the long-term, global impact of the deformation is studied. That is, the effect of severely straining and then unloading the system on the statistics on the yield stresses. Then, the short-term, local impact is investigated. Namely, how the local yield stress landscape evolves due to individual avalanches in the four different scenarios.
The local yield stress associated with a given sub-region of the sample (referred to as box in the rest of the paper) is assumed the be determined by the strength of the weakest of the activable substructures (links) within the box. (If the local yield stress is assigned to the whole simulation cell, we will simply refer to it as the yield stress or yield threshold). Thus, one may expect the local yield stress $\tau_{\mathrm{y}}$ to obey an extremal probability distribution, in particular, a first order statistics (i.e., minimum) distribution. Namely, if in the small strength limit the link strength $\tau_{\mathrm{link}}$ is powerlaw distributed as $F_{\text{link}}(\tau_{\text{link}})\propto\tau_{\text{link}}^{k}$, $F_{\text{link}}$ being the cumulative distribution function (CDF) of $\tau_{\mathrm{link}}$, then the emergent extremal distribution is of Weibull type [30, 31, 32, 33, 34] with a CDF
@@
Here $\lambda$ and $k$ are the so-called scale and shape parameters, respectively. The former is proportional to the mean value of the variable and the latter describes the asymptotic behavior at the $\tau_{\mathrm{y}}\rightarrow{}0$ limit. Here we consider $\tau^{\prime}_{\mathrm{y}}=\tau_{\mathrm{y}}-\tau_{\mathrm{init}}$ where $\tau_{\mathrm{init}}$ is the initial external stress which is zero for all scenarios except scenario III. The statistics of yield thresholds $\tau^{\prime}_{\mathrm{y}}$ (corresponding to the whole simulation cell) were obtained during loading the sample with a quasi-statically increasing spatially homogeneous external stress, $\tau_{\mathrm{y}}$ being the critical stress at which the first avalanche occurs. The event is detected based on thresholding: the avalanche is considered to have started when both the mean dislocation velocity and its derivative w.r.t time exceed a predefined value. As it is shown in Fig. 13 the statistics of the yield thresholds $\tau^{\prime}_{\mathrm{y}}$ indeed obeys the Weibull distribution with a shape parameter $k\approx 1.1\pm 0.1$ robust to the change of the extent $Q$ of quenched disorder and to the deformation history of the sample. The scale parameter $\lambda$, however, depends on $Q$ and the deformation history. The most significant effect is that prestraining (scenario IV) immensely increases the yield stress values due to the elimination of weak substructures that are easy to trigger during the loading stage.
@@
Above it was shown that while some features (e.g. the shape parameter) of the yield threshold distribution is robust, the straining history can affect the statistics of local yield stresses. Now let us focus on the short-term changes of local yield stresses (corresponding to smaller subsystems) after individual dislocation avalanches. To this end the following procedure is applied. Each configuration is divided into $8\times 8$ square shaped disjoint boxes. Then the local yield stress of each box is probed by quasi-statically increasing spatially homogeneous external stress while keeping the dislocations outside the box fixed. The computation of local yield stresses is done with the thresholding-based detection of avalanches described above. The local yield stresses are first computed for all boxes based on the first avalanches. Then, quasi-static global loading is applied to the relaxed configurations. That is, all dislocations are affected by the homogeneous external stress and all are mobile. The configurations are loaded until the first avalanche occurs (that is, until the yield threshold of the whole simulation cell). This, naturally results in the spatial redistribution of dislocations which leads to the change of the local yield stress of each fictitious box. The local yield stresses are measured after this global loading as well. The change of $\tau_{\mathrm{y}}$ is denoted with $\Delta\tau_{\mathrm{y}}$. For each configuration a map of $\Delta\tau_{\mathrm{y}}$ is created. The box with the largest change (in terms of absolute value) is considered the be the core of the avalanche was. Each map is rearranged by moving the core box into the center while keeping the relative positions (quantified by coordinates $\Delta x$ and $\Delta y$) of boxes (taking into account the PBC). These re-centered maps are than averaged (and symmetrized in order to decrease noise) over the ensembles of configurations of each of the four scenarios. The so obtained maps of $|\Delta\tau_{\mathrm{y}}|$ and $\Delta\tau_{\mathrm{y}}$ are shown in Figs. 14 and 15, respectively.
@@
The maps indicate that at the core box (that is, where the dislocation avalanche emerged) $\langle|\Delta\tau_{\mathrm{y}}|\rangle$ is significantly larger than in any other box and $\langle\Delta\tau_{\mathrm{y}}\rangle>0$, that is, an immense hardening is observed. This is not surprising since after the activation of the first avalanche the next activable link is probably significantly harder to trigger. This can be explained by the assumption (consistent with the numerically obtained Weibull distributions of $\tau_{\mathrm{y}}$) that the low-strength asymptotic distribution of link strengths obeys a power-law. This is the part of the distribution that matters during weakest link activation. Since at the low-strength tail of the distribution the probability density is low, independently of the redistribution of dislocations, one’s intuition would be that the second avalanche in this core box is significantly harder to trigger than the first one. The other boxes are also affected by the avalanche but to a smaller extent and typically softening can be observed instead of hardening. The probability density functions of $\Delta\tau_{\mathrm{y}}$ are shown in Fig. 16. The core boxes almost always (with more than 93% probability in each scenario) harden with $\Delta\tau_{\mathrm{y}}$ approximately obeying an exponential distribution. The non-core boxes typically soften (with more than 76% probability in each scenario) due to the increased external stress and the change of internal stresses resulted by the dislocation rearrangements.
@@
Figure 17 shows the dependence of $\langle|\Delta\tau_{\mathrm{y}}|\rangle$ on the distance $\Delta r=\sqrt{\Delta x+\Delta y}$ measured from the avalanche core. In every scenario the impact of the avalanche on the local yield stresses decay according to
@@
As in relaxed $Q=10$ systems (scenario II) the dynamical correlations are weaker and shorter-ranged than in their $Q=0$ counterparts (scenario I), not surprisingly, the same can be observed in terms of the magnitude of $\langle|\Delta\tau_{\mathrm{y}}|\rangle$ and the exponent $\beta$. More surprisingly, $\beta$ remain the same during severe straining (scenario III), although, the avalanche impact is stronger (comparable with scenario I) indicated by the increase in the magnitude of $\langle|\Delta\tau_{\mathrm{y}}|\rangle$. In the prestraining case (scenario IV) $\beta\approx 0$, that is, the softening is comparable in all boxes. This is the results of the first avalanche being triggered after an increment of external stress much larger than in any other scenarios which leads to a larger extent of rearrengement (in the form of drifting dislocation motion) prior to the avalanche. This is probably related to the elimination of substructures that are easy to trigger during the prestraining (which was also suggested in Ref. [35]).
@@
It is generally true (regardless of the scenario) that the boxes that are vertically or horizontally aligned with the core box are more impacted by the avalanche while the diagonally positioned boxes are less affected. This angular dependence can be related to the stress field induced by the rearrangement of the dislocations by the inspection of the Eshelby stress field of a displaced dislocation [36]. The shear stress field of a dislocation displaced with an infinitesimally small $\Delta x$ changes with $\Delta\tau=\partial_{x}\tau_{d}\Delta x=\frac{\cos(4\varphi)}{r^{2}}\Delta x$ where $\varphi$ is the angular polar coordinate and $\tau_{\mathrm{d}}$ is the shear stress field of a dislocation (see Eq. (1)). If it is assumed that (i) during an avalanche the displacement of dislocations is typically small, (ii) most of the dislocations move in the direction consistent with the external stress and (iii) the dislocation motion is the most intense in a relatively local and compact inner core of the avalanche, the change of the internal stress field has roughly a $\cos(4\varphi)$ type angular dependence (consistently with the results in Ref. [29]) which strengthens the effect of the external stress in the vertical and horizontal directions and weakens it diagonally. This leads to the observed enhanced softening vertically and horizontally and the mitigated softening diagonally.
@@
Figure 17: The average of the absolute value of the change of the local yield stresses $\langle|\Delta\tau_{\mathrm{y}}|\rangle$ in the boxes as a function of the distance $\Delta r$ of their middle from the middle of the box that contains the core of the avalanche (that is, where $|\Delta\tau_{\mathrm{y}}|$ is maximal). ‘$+$’: box horizontally or vertically aligned with the core box, ‘$\times$’: box diagonally positioned in comparison to the core box, $\circ$’: every other box. The inset shows the exponent $\beta$ characterizing the powerlaw decay with the distance from the core box. $L$ and $\tau_{0}$ are the linear size of the simulation cell and a constant defined in the main text, respectively.
@@ IV Conclusion and outlook
Besides the correlations, the yield stress (corresponding to whole configurations) and local yield stresses (corresponding to subsystems) were studied. The results showed that the yield stresses (consistently with the weakest-link principle) obey Weibull distribution with a shape parameter robust to the introduction of quenched disorder and to the deformation history. It was observable, however, that the loading and subsequent unloading of a configuration results in a remarkable hardening (that is, a shift of the scale parameter of the yield stress distribution). The analysis of local yield stresses indicate that the local hardness of the material changes significantly when an avalanche occurs due to the rearrangement of dislocations. At the core of the avalanche an immense hardening is observable. In the more remote vicinity, however, the material is (on average) softening and the impact is significantly weaker. The softening effect diminishes as the distance from the core of the avalanche is increased. More precisely, the mean change of the local yield stress decays according to powerlaw with an exponent dependent on the extent of quenched disorder and the straining history and it exhibits an angular dependence consistent with the Eshelby stress field.

### v2

@@ III.5 Yield stresses and local yield stresses
After studying the correlations and the avalanche behavior now we turn to one of the main questions if this work: how are the local yield stresses affected by the avalanches and the deformation of the model sample. Firstly, the long-term, global impact of the deformation is studied. That is, the effect of severely straining and then unloading the system on the statistics on the yield stresses. Then, the short-term, local impact is investigated. Namely, how the local yield stress landscape evolves due to individual avalanches in the case of the four different conditions.
The local yield stress associated with a given sub-region of the sample (referred to as box in the rest of the paper) is assumed to be determined by the strength of the weakest of the activable substructures (links) within the box. (If the local yield stress is assigned to the whole simulation cell, we will simply refer to it as the yield stress or yield threshold). Thus, one may expect the local yield stress $\tau_{\mathrm{y}}$ to obey an extremal probability distribution, in particular, a first order statistics (i.e., minimum) distribution. Namely, if in the small strength limit the link strength $\tau_{\mathrm{link}}$ is power-law distributed as $F_{\text{link}}(\tau_{\text{link}})\propto\tau_{\text{link}}^{k}$, $F_{\text{link}}$ being the cumulative distribution function (CDF) of $\tau_{\mathrm{link}}$, then the emergent extremal distribution is of Weibull type [34, 35, 36, 37, 38] with a CDF
@@
Here $\lambda$ and $k$ are the so-called scale and shape parameters, respectively. The former is proportional to the mean value of the variable and the latter describes the asymptotic behavior at the $\tau_{\mathrm{y}}\rightarrow{}0$ limit. Here we consider $\tau^{\prime}_{\mathrm{y}}=\tau_{\mathrm{y}}-\tau_{\mathrm{init}}$ where $\tau_{\mathrm{init}}$ is the initial external stress which is zero for all conditions except condition III. The statistics of yield thresholds $\tau^{\prime}_{\mathrm{y}}$ (corresponding to the whole simulation cell) were obtained during loading the sample with a quasi-statically increasing spatially homogeneous external stress, $\tau_{\mathrm{y}}$ being the critical stress at which the first avalanche occurs. The event is detected based on thresholding: the avalanche is considered to have started when both the mean dislocation velocity and its derivative w.r.t time exceed a predefined value. As it is shown in Fig. 13 the statistics of the yield thresholds $\tau^{\prime}_{\mathrm{y}}$ indeed obeys the Weibull distribution with a shape parameter $k\approx 1.1\pm 0.1$ robust to the change of the extent $Q$ of quenched disorder and to the deformation history of the sample. The scale parameter $\lambda$, however, depends on $Q$ and the deformation history. The most significant effect is that prestraining (condition IV) immensely increases the yield stress values due to the elimination of weak substructures that are easy to trigger during the loading stage.
@@
Above it was shown that while some features (e.g. the shape parameter) of the yield threshold distribution is robust, the straining history can affect the statistics of local yield stresses. Now let us focus on the short-term changes of local yield stresses (corresponding to smaller subsystems) after individual dislocation avalanches. To this end the following procedure is applied. Each configuration is divided into $8\times 8$ square shaped disjoint boxes. Then the local yield stress of each box is probed by quasi-statically increasing spatially homogeneous external stress while keeping the dislocations outside the box fixed. The computation of local yield stresses is done with the thresholding-based detection of avalanches described above. The local yield stresses are first computed for all boxes based on the first avalanches. Then, quasi-static global loading is applied to the relaxed configurations. That is, all dislocations are affected by the homogeneous external stress and all are mobile. The configurations are loaded until the first avalanche occurs (that is, until the yield threshold of the whole simulation cell). This, naturally results in the spatial redistribution of dislocations which leads to the change of the local yield stress of each fictitious box. The local yield stresses are measured after this global loading as well. The change of $\tau_{\mathrm{y}}$ is denoted with $\Delta\tau_{\mathrm{y}}$. For each configuration a map of $\Delta\tau_{\mathrm{y}}$ is created. The box with the largest change (in terms of absolute value) is considered to be the core of the avalanche was. Each map is rearranged by moving the core box into the center while keeping the relative positions (quantified by coordinates $\Delta x$ and $\Delta y$) of boxes (taking into account the PBC). These re-centered maps are than averaged (and symmetrized in order to decrease noise) over the ensembles of configurations of each of the four conditions. The so obtained maps of $|\Delta\tau_{\mathrm{y}}|$ and $\Delta\tau_{\mathrm{y}}$ are shown in Figs. 14 and 15, respectively.
@@
The maps indicate that at the core box (that is, where the dislocation avalanche emerged) $\langle|\Delta\tau_{\mathrm{y}}|\rangle$ is significantly larger than in any other box and $\langle\Delta\tau_{\mathrm{y}}\rangle>0$, that is, an immense hardening is observed. This is not surprising since after the activation of the first avalanche the next activable link is probably significantly harder to trigger. This can be explained by the assumption (consistent with the numerically obtained Weibull distributions of $\tau_{\mathrm{y}}$) that the low-strength asymptotic distribution of link strengths obeys a power-law. This is the part of the distribution that matters during weakest link activation. Since at the low-strength tail of the distribution the probability density is low, independently of the redistribution of dislocations, one’s intuition would be that the second avalanche in this core box is significantly harder to trigger than the first one. The other boxes are also affected by the avalanche but to a smaller extent and typically softening can be observed instead of hardening. The probability density functions of $\Delta\tau_{\mathrm{y}}$ are shown in Fig. 16. The core boxes almost always (with more than 93% probability in each condition) harden with $\Delta\tau_{\mathrm{y}}$ approximately obeying an exponential distribution. The non-core boxes typically soften (with more than 76% probability in each condition) due to the increased external stress and the change of internal stresses resulted by the dislocation rearrangements.
@@
Figure 17 shows the dependence of $\langle|\Delta\tau_{\mathrm{y}}|\rangle$ on the distance $\Delta r=\sqrt{\Delta x+\Delta y}$ measured from the avalanche core. In every condition the impact of the avalanche on the local yield stresses decay according to
@@
As in relaxed $Q=10$ systems (condition II) the dynamical correlations are weaker and shorter-ranged than in their $Q=0$ counterparts (condition I), not surprisingly, the same can be observed in terms of the magnitude of $\langle|\Delta\tau_{\mathrm{y}}|\rangle$ and the exponent $\beta$. More surprisingly, $\beta$ remain the same during severe straining (condition III), although, the avalanche impact is stronger (comparable with condition I) indicated by the increase in the magnitude of $\langle|\Delta\tau_{\mathrm{y}}|\rangle$. In the prestraining case (condition IV) $\beta\approx 0$, that is, the softening is comparable in all boxes. This is the results of the first avalanche being triggered after an increment of external stress much larger than in the case of any other conditions which leads to a larger extent of rearrangement (in the form of drifting dislocation motion) prior to the avalanche. This is probably related to the elimination of substructures that are easy to trigger during the prestraining (which was also suggested in Ref. [39]).
@@
It is generally true (regardless of the condition) that the boxes that are vertically or horizontally aligned with the core box are more impacted by the avalanche while the diagonally positioned boxes are less affected. This angular dependence can be related to the stress field induced by the rearrangement of the dislocations by the inspection of the Eshelby stress field of a displaced dislocation [40]. The shear stress field of a dislocation displaced with an infinitesimally small $\Delta x$ changes with $\Delta\tau=\partial_{x}\tau_{d}\Delta x=\frac{\cos(4\varphi)}{r^{2}}\Delta x$ where $\varphi$ is the angular polar coordinate and $\tau_{\mathrm{d}}$ is the shear stress field of a dislocation (see Eq. (1)). If it is assumed that (i) during an avalanche the displacement of dislocations is typically small, (ii) most of the dislocations move in the direction consistent with the external stress and (iii) the dislocation motion is the most intense in a relatively local and compact inner core of the avalanche, the change of the internal stress field has roughly a $\cos(4\varphi)$ type angular dependence (consistently with the results in Ref. [33]) which strengthens the effect of the external stress in the vertical and horizontal directions and weakens it diagonally. This leads to the observed enhanced softening vertically and horizontally and the mitigated softening diagonally.
We note that the observations above are not limited to the box size corresponding to the resolution of $8\times 8$. Our prior investigation revealed that for finer resolutions than $8\times 8$ empty boxes may appear which formally have infinite local yield stress in our model. Therefore, the resolution of $8\times 8$ was chosen since it is suitable for our analysis and provides a more detailed picture than coarser resolutions. However, it should be noted that our analyses (not shown here) conducted at resolutions $4\times 4$ and $6\times 6$ corroborates that the general conclusions drawn above (e.g., the $\cos(4\varphi)$-type anisotropy, the approximately power-law decay of $|\Delta\tau_{\mathrm{y}}|$ and the differences between the four conditions) do not depend on the actual box size.
@@
Figure 17: The average of the absolute value of the change of the local yield stresses $\langle|\Delta\tau_{\mathrm{y}}|\rangle$ in the boxes as a function of the distance $\Delta r$ of their middle from the middle of the box that contains the core of the avalanche (that is, where $|\Delta\tau_{\mathrm{y}}|$ is maximal). ‘$+$’: box horizontally or vertically aligned with the core box, ‘$\times$’: box diagonally positioned in comparison to the core box, $\circ$’: every other box. The inset shows the exponent $\beta$ characterizing the power-law decay with the distance from the core box. $L$ and $\tau_{0}$ are the linear size of the simulation cell and a constant defined in the main text, respectively.
@@ IV Conclusion and outlook
Besides the correlations, the yield stress (corresponding to whole configurations) and local yield stresses (corresponding to subsystems) were studied. The results showed that the yield stresses (consistently with the weakest-link principle) obey Weibull distribution with a shape parameter robust to the introduction of quenched disorder and to the deformation history. It was observable, however, that the loading and subsequent unloading of a configuration results in a remarkable hardening (that is, a shift of the scale parameter of the yield stress distribution). The analysis of local yield stresses indicate that the local hardness of the material changes significantly when an avalanche occurs due to the rearrangement of dislocations. At the core of the avalanche an immense hardening is observable. In the more remote vicinity, however, the material is (on average) softening and the impact is significantly weaker. The softening effect diminishes as the distance from the core of the avalanche is increased. More precisely, the mean change of the local yield stress decays according to power-law with an exponent dependent on the extent of quenched disorder and the straining history and it exhibits an angular dependence consistent with the Eshelby stress field.
@@
Acknowledgements.
Support by the National Research, Development and Innovation Fund of Hungary (contract numbers: NKFIH-FK-138975) is acknowledged. Supported by the ÚNKP-21-1 New National Excellence Program of the Ministry for Innovation and Technology from the source of the National Research, Development and Innovation Fund.